unseen:

The sum of eight and a number is one hundred. What is the number? 12.5 88 92 108

unseen:

d

DanielleLove:

8+a number=100

DanielleLove:

100-8=92

unseen:

so it would be D

DanielleLove:

no

unseen:

because your adding not subtracting

DanielleLove:

ok 8+x=100 to find x work pemdas backwards 100-8=92

unseen:

looks like i was correct

DanielleLove:

or you could just be trolling because I'm 100% sure of my answer

unseen:

well im not so back off

unseen:

and i closed it so good bye